# ArtWhisperer: A Dataset for Characterizing Human-AI Interactions in Artistic Creations

### v1

@@ Abstract
As generative AI becomes more prevalent, it is important to study how human users interact with such models. In this work, we investigate how people use text-to-image models to generate desired target images. To study this interaction, we created ArtWhisperer, an online game where users are given a target image and are tasked with iteratively finding a prompt that creates a similar-looking image as the target. Through this game, we recorded over 50,000 human-AI interactions; each interaction corresponds to one text prompt created by a user and the corresponding generated image. The majority of these are repeated interactions where a user iterates to find the best prompt for their target image, making this a unique sequential dataset for studying human-AI collaborations. In an initial analysis of this dataset, we identify several characteristics of prompt interactions and user strategies. People submit diverse prompts and are able to discover a variety of text descriptions that generate similar images. Interestingly, prompt diversity does not decrease as users find better prompts. We further propose to a new metric the study the steerability of AI using our dataset. We define steerability as the expected number of interactions required to adequately complete a task. We estimate this value by fitting a Markov chain for each target task and calculating the expected time to reach an adequate score in the Markov chain. We quantify and compare AI steerability across different types of target images and two different models, finding that images of cities and natural world images are more steerable than artistic and fantasy images. These findings provide insights into human-AI interaction behavior, present a concrete method of assessing AI steerability, and demonstrate the general utility of the ArtWhisperer dataset.
@@ 1 Introduction
Direct human interaction with AI models has become widespread following the public release of text-to-text models like GPT-4 [17] and PaLM 2 [1] and text-to-image models like Stable Diffusion [24]. These models have seen rapid interest and adoption in diverse industries including engineering, creative writing, art, education, medicine, and law [6, 16, 9, 4, 22, 3, 28]. A large reason for why this rapid adoption has been possible is that people, with no understanding of how the model works or its limitations, are able to interact with and steer the model to complete a wide variety of tasks.
@@
One of the key challenges for developing these models is aligning their output to human-written input text. This is made especially challenging by the broad domain of use cases as well as the diverse prompting styles of different users. Many approaches can be categorized into the broad label of “prompt engineering” where specific strategies for prompting are used to steer a model [18, 13, 37, 30, 31]. Great success has also been found by fine-tuning models with relatively small datasets to better align a model to follow human instructions [19], to condition the model to respond in a specific style [8], or behave differently to specified prompts [36, 7].
@@
In this work, we take special interest in the fact that human interaction with these models often is an iterative process and develop a dataset to help investigate how people interact with these models with a goal of improving human-AI interaction. To study this interaction, we created an interactive game where players try to find an optimal prompt for a given task (see Figure 1). In particular, we focus on text-to-image models and ask the player to generate a similar image (AI Image) to a given target image. The player is allowed to iterate on their prompt, using the previously generated image(s) as feedback to help them adjust their prompt. A score is also provided as feedback to help the user calibrate how “close” they are to a similar image.
@@
Using this setup, we collected interaction data on 51,026 interactions from 2,250 players across 191 unique target images. The target images were selected from a diverse set of AI-generated and natural images. We also collected a separate dataset of 4,572 interactions, 140 users, and 51 unique target images in a more controlled setting with two different diffusion models to assess the robustness of our findings.
@@
Based on this data, we find several interesting patterns in how people interact with AI models. Players are able to find many different ways to achieve high scores; in particular, they find a diverse set of prompts that all result in a player-generated image similar to the target image. We also find that players tend to make small, iterative updates to their prompts as they steer the AI, with each update improving their image with a moderate success rate ($40-60\%$ for most target images). Based on these findings, we define and evaluate a metric model steerability based on the stopping time of an empirical Markov model. We use this metric to assess steerability across different groups of images and across two AI models.
Ethical considerations
One of the main goals of this work is to help improve the quality of human-AI interaction. our dataset and findings provide quantitative insights on how people interact with generative AI and can potentially be used to design AI that are easier for people to use. It does not address the broader concern that bad actors may abuse generative AI models.
Our contributions
@@
We release a public dataset on human interaction with an AI model. To our knowledge, this is the first such dataset showing repeated interactions of people with a text-to-image model to accomplish specified tasks. We also provide an initial analysis of this data and propose a simple-to-calculate metric for assessing model steerability. Our dataset and associated code is made available at https://github.com/kailas-v/ArtWhisperer.
@@
Figure 1: Interface of ArtWhisperer, our human-AI interaction game. Prompts entered on right. Target image and player-generated image on left. In the lower right is a section where previous prompts and scores are displayed.
Related Works
@@
Datasets for human interaction with text-to-text or text-to-image models typically focus on single interactions (a single user prompt and the resulting model output) and generally do not provide users with a specific task. Public text-to-image interaction datasets typically contain the generated AI images or prompts [26], paired data with both images and prompts [29], and data also paired with some form of human preference rating [20, 34, 35, 11]. These datasets generally rely on scraping online repositories like Lexica [12] or Discord servers focused on AI Art generation. Though some of these datasets include metadata like timestamp and a user ID allowing reconstruction of prompt iteration, there is no guarantee the user has the same target image in mind or whether the user has changed their desired output over time. Public text-to-text datasets are much more limited as the best performing models are generally accessible only through APIs with no public repositories that aggregate user interactions. While some researchers have investigated how human-AI interaction for text-to-text can be improved through various tools [32, 33], the amount data collected is limited and not publicly available. There are also repositories containing prompt strategies for various tasks [2], but no human interaction component.
@@
We seek to rectify two of the shortcomings of the existing datasets–namely, that they do not contain extended interactions as the user attempts to steer the AI, and they do not have a predefined goal. In our work, we create a controlled environment where we allow extended interactions and have a known goal for human users. As shown by our initial analysis, our dataset may enable deeper understanding of user prompting strategies and assessing model steerability.
@@ 2 Interaction Game
In the game, players are shown a target image. A few example target images are provided in Figures 24. Players are also provided a limited interface to a text-to-image model; in particular, they are given access to the “positive prompt” and “negative prompt” inputs to the Stable Diffusion (SD) v2.1 model [25]. Here, the positive prompt should contain what the user would like to see in the generated image, while the negative prompt should contain what the user would like omitted from the generated image. Upon inputting a prompt (which often only includes a positive prompt and no negative prompt), the player is shown the image generated by the AI model, along with a score based on similarity between their generated image and the target image. This interface is shown in Figure 1.
Note that players only have access to the model through the positive and negative prompts. In particular, all other parameters are fixed, including the random seed. We opted to fix the random seed for each target image so that all players will generate the same image given the same prompt.
@@ 2.1 How Target Images are Selected
We randomly sample target images from two sources. The first is a collection of Wikipedia pages, and the second is a dataset of prompts AI artists have used with SD [26]. In addition to sampling target images, we need to ensure the task is feasible to users. As we do not allow users to adjust the seed or other parameters of the model, we need to ensure the selected model parameters can generate reasonably similar images to the target image. We find that selecting an appropriate random seed is sufficient, and fix all other model parameters (see Section 2.3 for details). Below, we describe how we randomly sample target images and select the random seed.
@@
A collection of 35 Wikipedia pages on various topics including art, nature, cities, and various people. A full list of pages sampled from is provided in the Appendix. From these pages, we scraped 670 figures licensed under the Creative Commons license. These figures were then filtered by which had captions, as well as which images were JPG or PNG images (i.e., not animated, and not PDF files), resulting in 557 images.
@@
$i^{*}:=\min_{i=1,\dots,50}\frac{||C(x_{i})-C(t_{i})||_{2}}{||C(x_{i})||_{2}\cdot||C(t_{i})||_{2}}.$
@@
A collection of 2,000 AI-art prompts are randomly sampled from the Stable Diffusion Prompts dataset [26]. For each prompt, $p_{i}^{*}$, we generate two sets of images using 60 unique random seeds: the first set, $S_{i,1}=\{(x_{i},s_{i}):i=1,\dots,10\}$ and $S_{i,2}=\{(x_{i},s_{i}):i=1,\dots,50\}$. We select the target image, $t_{i_{1}^{*}}$, from $S_{i,1}$:
$i_{1}^{*}:=\min_{i=1,\dots,10}\text{median}\left(\left\{\frac{||C(x_{i,1})-C(x_{j,2})||_{2}}{||C(x_{i,1})||_{2}\cdot||C(x_{j,2})||}:j=1,\dots,50\right\}\right)$
@@
$i_{2}^{*}:=\min_{i=1,\dots,50}\frac{||C(x_{i,2})-C(t_{i_{1}^{*}})||_{2}}{||C(x_{i,2})||_{2}\cdot||C(t_{i_{1}^{*}})||_{2}}.$
The intuition here is that $t_{i_{1}^{*}}$ is more representative of the types of images we may expect given the fixed prompt, $p_{i}^{*}$.
@@ 2.2 Scoring Function
$score(x_{i},t_{i})=\max(0,\min(100,\alpha\cdot\frac{\langle CLIP(x_{i}),CLIP(t_{i})\rangle}{||CLIP(x_{i})||_{2}\cdot||CLIP(t_{i})||_{2}}+\beta)),$
for generated image $x_{i}$, target image $t_{i}$, and constants $\alpha,\beta$. Note the range of $score(x_{i},t_{i})$ is integers in the interval $[0,100]$. Details on how $\alpha,\beta$ are selected parameters are provided in the Appendix.
While this scoring function is often reasonable, it does not always align with the opinions of a human user. To assess how well $score(x_{i},t_{i})$ follows a user’s preferences, we acquire ratings from a subset of users (see ArtWhisperer-Validation in Section 2.4). We find $\text{score}(x_{i},t_{i})$ has a Pearson correlation coefficient of 0.579 indicating reasonable agreement. Further assessment is performed in Section 4.3.
2.3 Additional Technical Details
For the generative model, we use SD v2.1 [25]. We use the DPM Multi-step Scheduler [14], and run the model for 20 iterations. AI-generated target images use the same parameters, but we run the model for 50 iterations. The primary reason for this difference was to limit latency for players; running for 20 iterations resulted in a latency of 1-3 seconds depending on the player’s internet connection. All images are generated at size 512 $\times$ 512.
Figure 2: Example user trajectories. In each row, the first 4 images show the prompt progression for a given user. Target image shown in green column. Plot shows the average score trajectory across all users for this target image (blue) and the specific user’s full score trajectory (red). Orange circles indicate the displayed images.
@@ 2.4 Dataset Overview
We collected two datasets as part of this work, which we term ArtWhisperer and ArtWhisperer-Validation. We use ArtWhisperer for most analysis and results; for some of the results in Sections 2.2, 4.2,and 4.3, we also use ArtWhisperer-Validation.
Data was collected from March-May 2023. IRB approval was obtained.
ArtWhisperer: A public version of our game was released online at https://artwhisperer.io/ and we collected data from consenting users playing the game. Three new images were released each day, and users can play during as many days as they would like. Users were anonymous and we only collected data related to the prompts submitted to ensure privacy of users. While we expect some users played the game across multiple days, we did not track them and so do not know the true number of unique users across days. A summary of the ArtWhisperer dataset is provided in Table 1. In total, we have 2,250 players corresponding to 51,026 interactions across 191 target images. We further break down the data across several common categories the images relate to. In Figure 3, we show a density plot of how many queries are submitted by players across different target images.
ArtWhisperer-Validation: A second version of the game (with a near identical interface) was released to paid crowd workers on Prolific [21]. The crowd workers were compensated at a rate of \$$\$12.00$ per hour for roughly 20 minutes of their time. Workers played the game across 5 randomly selected target images from a pre-selected subset of 51 target images chosen to have diverse content. Workers were also asked to rate each of their images on a scale of 1-10 (i.e., self-scoring their generated images). In total, we collected data on 4,572 interactions, corresponding to 140 users and 51 unique target images across two different diffusion models, SD v2.1 and SD v1.5. Additional details on the ArtWhisperer-Validation dataset are provided in the Appendix.
Table 1: ArtWhisperer Dataset Overview. Each row contains summary data for a different subset of the dataset. Subsets may overlap. Similar information for ArtWhisperer-Validation is in the Appendix.
@@
Figure 3: Left, Distribution of # of queries per target image. The average number of queries per image is 9.18. Right, Distribution of the # of words submitted in a query. The average number of words submitted in a positive and negative prompt are 20.02 and 2.32 respectively.
@@ 3 Prompt Diversity
Across images, regardless of image content, people submit a diverse set of prompts. Notably, people are able to achieve high scores with a diverse set of prompts. While not necessarily surprising that there are multiple ways to achieve a good score (this indicates that the score in our game has multiple local maxima for any given target image), it is interesting that the high-scoring prompts submitted by users remains diverse. Some examples are shown in Figure 4.
We quantify prompt diversity by looking at the distribution of prompts in the text embedding space. In particular, we use the CLIP text embedding [23], and measure the distance in the embedding space between a given prompt and the average submitted prompt for the corresponding target image. Though the results presented here are based on the CLIP embedding, we do find similar results for other standard text embeddings (see Supplement).
@@ 3.1 Diverse prompts used for high scores
In the left of Figure 5, we plot the distribution of distances between the first prompt (in blue) and the last prompt (in orange) to the average prompt for the corresponding target image. Despite the average score improving from 51.9 to 70.3 (out of 100) indicating a significant improvement in score, prompt diversity does not significantly diminish. That is, users do not converge to similar prompts to achieve high scores. Similar analysis of the image embedding space suggests image diversity decreases (see Appendix for details). This makes sense given that the score is linearly related to distance in the CLIP image embedding space. This analysis suggests that, over the course of interaction with the AI model, the distribution of user prompts undergoes a transformation (to be closer to the target image), but does not change in diversity.

### v4

@@ Abstract
In this work, we investigate how people use text-to-image models to generate desired target images. To study this interaction, we created ArtWhisperer, an online game where users are given a target image and are tasked with iteratively finding a prompt that creates a similar-looking image as the target. Through this game, we recorded over 50,000 human-AI interactions; each interaction corresponds to one text prompt created by a user and the corresponding generated image. The majority of these are repeated interactions where a user iterates to find the best prompt for their target image, making this a unique sequential dataset for studying human-AI collaborations. In an initial analysis of this dataset, we identify several characteristics of prompt interactions and user strategies. People submit diverse prompts and are able to discover a variety of text descriptions that generate similar images. Interestingly, prompt diversity does not decrease as users find better prompts. We further propose a new metric to quantify AI model steerability using our dataset. We define steerability as the expected number of interactions required to adequately complete a task. We estimate this value by fitting a Markov chain for each target task and calculating the expected time to reach an adequate score. We quantify and compare AI steerability across different types of target images and two different models, finding that images of cities and nature are more steerable than artistic and fantasy images. We also evaluate popular vision-language models to assess their image understanding and ability to incorporate feedback. These findings provide insights into human-AI interaction behavior, present a concrete method of assessing AI steerability, and demonstrate the general utility of the ArtWhisperer dataset.
@@ 1 Introduction
Direct human interaction with AI models has become widespread following a number of technical innovations improving the quality of text-to-text [5, 30, 1] and text-to-image models [37, 36], enabling the public release of high-quality AI-based services like ChatGPT [8], Bard [4], and Midjourney [25]. These models have seen rapid interest and adoption largely due to the ability of the general public to interact with and steer the AI in diverse contexts including engineering, creative writing, art, education, medicine, and law [10, 27, 16, 7, 34, 6, 42].
@@
A key challenge in developing these models is aligning their output to human inputs. This is made challenging by the broad domain of use cases as well as the diverse prompting styles of different users. Many approaches can be categorized as “prompt engineering,” where specific strategies for prompting are used to steer a model [29, 22, 53, 45, 46]. Great success has also been found by fine-tuning models using relatively small datasets to follow human instructions [30], respond in a specific style [15], or behave differently to specified prompts [52, 12].
@@
In this work, we take interest in the fact that human interaction with these models is often an iterative process. We develop a dataset to study this interaction. The dataset is collected through an interactive game we created where players try to find an optimal prompt for a given task (see Figure 1). In particular, we focus on text-to-image models and ask the player to generate a similar image (AI Image) to a given target image. The player is allowed to iterate on their prompt, using the previously generated image(s) as feedback to help them adjust their prompt. A score is also provided as feedback to help the user calibrate how “close” they are to a similar image.
@@
Using this setup, we collected data on 51,026 interactions from 2,250 players across 191 unique target images. The target images were selected from a diverse set of AI-generated and natural images. We also collected a separate dataset of 4,572 interactions, 140 users, and 51 unique target images in a more controlled setting to assess the robustness of our findings.
@@
Based on this data, we find several interesting patterns in how people interact with AI models. Players discover a diverse set of prompts that all result in images similar to the target. To discover these prompts, players typically make small, iterative updates to their prompts. Each update improves their image with a moderate success rate ($40-60\%$ for most target images). Based on these findings, we define and evaluate a metric for model steerability using the stopping time of an empirical Markov model. We use this metric to assess steerability across image categories and across two AI models.
@@
Our contributions We release a public dataset on human interactions with an AI model. To our knowledge, this is the first such dataset showing repeated interactions of people with a text-to-image model to accomplish specified tasks. We also provide an initial analysis of this data and propose a simple-to-calculate metric for assessing model steerability. Additionally, we use our dataset and this steerability metric to evaluate the ability of vision-language models to utilize feedback. Our dataset and associated code is made available at https://github.com/kailas-v/ArtWhisperer.
@@
Figure 1: Interface of the ArtWhisperer game. Prompts entered on right. Target (goal) image and player-generated image on left. Previous prompts and scores are displayed in the lower right.
@@
Related Works Human-AI interaction datasets for text-to-text and text-to-image models typically focus on single interactions and generally do not provide users with a specific task. Public text-to-image interaction datasets typically contain the generated AI images and/or prompts [39, 44] and optionally some form of human preference rating [32, 49, 50, 19]. These datasets generally rely on scraping online repositories like Lexica [20] or Discord servers focused on AI art. Though some of these datasets include metadata that may allow for reconstruction of prompt iteration, there is no guarantee the user has the same desired output in mind over the iteration. Public text-to-text interaction datasets are much more limited as the best performing models are generally accessible only through APIs with no public user interaction datasets. While some researchers have investigated how human-AI interaction for text-to-text can be improved through various tools [47, 48], the amount of data collected is limited and not publicly available. There are also repositories containing prompt strategies for various tasks [3] but no human interaction component.
@@
We seek to rectify two of the shortcomings of the existing datasets–namely, that they do not contain extended interactions as the user attempts to steer the AI, and they do not have a predefined goal. In our work, we maintain a controlled environment for human users where we allow extended interactions towards a fixed goal. As shown by our initial analysis, our dataset may enable deeper understanding of user prompting strategies, assessing model steerability, and evaluating vision-language AI models.
@@ 2 Interaction Game
In our game, players are shown a target image (see Figures 2 and 4). Players are also given a limited interface to a text-to-image model, Stable Diffusion (SD) v2.1 model [38]. In particular, players can enter a “positive prompt” (describes the desired content) and a “negative prompt” (describes what should be omitted) to steer the AI model. All models hyperparameters including the seed are fixed. Upon inputting a prompt, the player is shown the image generated by the AI model, along with a similarity score between their generated image and the target image. The interface is shown in Figure 1.
@@ 2.1 How Target Images are Selected
We randomly sample target images from two sources. The first is a collection of Wikipedia pages, and the second is a dataset of prompts AI artists have used with SD [39]. In addition to sampling target images, we need to ensure the task is feasible to users. We do not allow users to adjust the seed or other parameters of the model, so we need to ensure the selected model parameters can generate reasonably similar images to the target image. We find that selecting an appropriate random seed is sufficient, and fix all other model parameters (see Appendix A.3 for examples of generated images and Appendix A.5 for details and discussion).
@@
A collection of 35 Wikipedia pages on various topics including art, nature, cities, and various people. A full list of pages sampled from is provided in Appendix 2. From these pages, we scraped 670 figures licensed under the Creative Commons license. These figures were then filtered by which had captions, as well as which images were JPG or PNG images (i.e., not animated, and not PDF files), resulting in 557 images.
@@
$i^{*}:=\min_{i=1,\dots,50}\left|\left|\frac{C(x_{i})}{||C(x_{i})||_{2}}-\frac{C(t_{i})}{||C(t_{i})||_{2}}\right|\right|_{2}$ (1)
Here, $s_{i^{*}}$ is selected to minimize the distance to the target image given the target prompt.
@@
A collection of 2,000 AI-art prompts are randomly sampled from the Stable Diffusion Prompts dataset [39]. For each prompt, $p_{i}^{*}$, we generate two sets of images. As before, we use 50 unique random seeds to select the seed, $s_{i^{*}}$ and an additional 10 random seeds to use for selecting the generated target image (so in total, we use 60 unique random seeds): the first set, $S_{i,1}=\{(x_{i,1},s_{i}):i=1,\dots,10\}$ and $S_{i,2}=\{(x_{i,2},s_{i}):i=1,\dots,50\}$. We select the target image, $t_{i_{1}^{*}}$, from $S_{i,1}$:
$i_{1}^{*}:=\min_{i=1,\dots,10}\text{median}$ ($\left(\left\{\left|\left|\frac{C(x_{i,1})}{||C(x_{i,1})||_{2}}-\frac{C(x_{j,2})}{||C(x_{j,2})||}\right|\right|_{2}:j=1,\dots,50\right\}\right)$ (2)
@@
$i_{2}^{*}:=\min_{i=1,\dots,50}\left|\left|\frac{C(x_{i,2})}{||C(x_{i,2})||_{2}}-\frac{C(t_{i_{1}^{*}})}{||C(t_{i_{1}^{*}})||_{2}}\right|\right|_{2}$ (3)
Here, $t_{i_{1}^{*}}$ is chosen to be more representative of the types of images we may expect given the fixed prompt, $p_{i}^{*}$. This is because $t_{i_{1}^{*}}$ is selected to be close to the center of the sampled images, $S_{i,2}$. The intuition for selecting $s_{i_{2}^{*}}$ is the same as selecting $s_{i^{*}}$ for the Wikipedia images.
@@ 2.2 Scoring Function
$score(x,t)=\max(0,\min(100,alpha\cdot\left|\left|\frac{C(x)}{||C(x)||_{2}}-\frac{C(t)}{||C(t)||_{2}}\right|\right|_{2}+\beta))$ (4)
for generated image $x$, target image $t$, and constants $\alpha,\beta$. Note the range of $score(x,t)$ is integers in the interval $[0,100]$. Details on how $\alpha,\beta$ are selected parameters are provided in Appendix A.4.
While this scoring function is often reasonable, it does not always align with the opinions of a human user. To assess how well $score(x,t)$ follows a user’s preferences, we acquire ratings from a subset of users (see ArtWhisperer-Validation in Section 2.3). We find $\text{score}(x,t)$ has a Pearson correlation coefficient of 0.579 indicating reasonable agreement. Further assessment is performed in Section 4.3 and discussed at length in Appendix A.14.
Figure 2: Example user trajectories. In each row, we show (1) a given user’s prompts, (2) the target image (rightmost image), and (3) a plot of this target image’s average score trajectory across users (blue), this user’s full score trajectory (red), and the displayed images (orange).
@@ 2.3 Dataset Overview
We collected two datasets: ArtWhisperer and ArtWhisperer-Validation. We use ArtWhisperer for most analysis and results; for some of the results in Sections 2.2, 4.2,and 4.3, we also use ArtWhisperer-Validation (when referenced). Data was collected from March-May 2023. IRB approval was obtained.
ArtWhisperer: A public version of our game was released online, with three new target images released daily. We collected data from consenting users playing the game. Users were not paid. Users were anonymous and we only collected data related to the prompts submitted to ensure privacy of users. While we expect some users played the game across multiple days, we did not track them. A summary of the ArtWhisperer dataset is provided in Table 1. In total, we have 2,250 (potentially non-unique) players corresponding to 51,026 interactions across 191 target images. Players interacted with the model SD v2.1. In Figure 3, we plot the number of queries submitted by players across different target images.
ArtWhisperer-Validation: The game (with a near identical interface) was also released as a controlled user study to paid crowd workers on Prolific [33]. The crowd workers were compensated at a rate of \$$\$12.00$ per hour for roughly 20 minutes of their time. Workers played the game across 5 randomly selected target images from a pre-selected subset of 51 target images chosen to have diverse content. Workers were also asked to rate each of their images on a scale of 1-10 (i.e., self-scoring their generated images). In total, we collected data on 4,572 interactions, corresponding to 140 users and 51 unique target images across two different diffusion models, SD v2.1 and SD v1.5. Additional details and demographic information are provided in Appendix A.7.
Table 1: ArtWhisperer Dataset Overview. Each row contains summary data for a different subset of the dataset. Subsets may overlap. Similar information for ArtWhisperer-Validation is in Appendix A.7.
@@
Figure 3: Left: Distribution of # of user queries per target image (average queries per image is 9.18). Right: Distribution of the # of words submitted in a query (average words submitted is 20.02 and 2.32 for positive and negative prompts respectively).
@@ 3 Prompt Diversity
We quantify prompt diversity by looking at the distribution of prompts in the text embedding space. In particular, we use the CLIP text embedding [35], though we do find the choice of embedding is not particularly important for our results (see Appendix A.8).
@@ 3.1 Diverse prompts used for high scores
First, we find that users do not converge in their prompt design but rather achieve similarly high scores through a diverse set of prompts. This result confirms the need for models to align to a wide range of user writing styles and descriptive techniques, and suggests a tradeoff between alignment to multiple users and specificity of a prompt (e.g., to support more users, individual prompts will necessarily have less descriptive ability). Examples of such diverse prompts are shown in Figure 4.
We quantify this finding in Figure 5, where we plot two metrics defined as follows. Let $z_{0},z_{n}$ be normalized embeddings of the initial and best prompt/image found by a user. Let $z^{*}$ be the normalized embedding of the target prompt/image. We define the difference in embedding distance to ground truth as $||z_{n}-z^{*}||_{2}-||z_{0}-z^{*}||_{2}$. In blue, we use the CLIP text embeddings of the prompts; in orange, we use the CLIP image embeddings of the generated images. We note two findings here: (1) the metric applied to the image embeddings is guaranteed to be non-positive as the embedding distance is monotonically decreasing with the score, and (2) the metric applied to the text embeddings is apparently symmetric around 00, indicating that unlike the image embedding, distance in the text embedding space does not monotonically decrease with score. Together, these findings illustrate that users tend to discover diverse prompts and do not converge in their prompt design.
Additionally, we find the distribution of prompts does not converge. In the left of Figure 6, we plot the distribution of distances between the first prompt (in blue) and the last prompt (in orange) to the average prompt for the corresponding target image. Despite the average score improving from 51.9 to 70.3 (out of 100) indicating a significant improvement in score, prompt diversity does not significantly diminish. That is, users do not converge to similar prompts to achieve high scores. Similar analysis of the image embedding space suggests image diversity decreases (Figure 5).